# Deep Insights into Cognitive Decline: A Survey of Leveraging Non-Intrusive Modalities with Deep Learning Techniques

### v1

@@ VII Discussion
Cognitive disorders. In addition to the natural decline associated with aging, several disorders can affect cognitive abilities. Depending on the disorder, the impact on cognitive functions can vary, and even in some cases, it can lead to dependency. Early detection is crucial to start professional treatment that can significantly improve a patient’s quality of life. These cognitive disorders include dementia, MCI, aphasia, Parkinson’s disease, and apathy. As mentioned earlier, our society is experiencing an aging trend, leading to an increasingly older population. This demographic shift results in a growing prevalence of age-related disorders, such as dementia. Most research in this field focuses on dementia, primarily due to the availability of relevant datasets. However, recent studies have also begun addressing other conditions, such as MCI, and in some cases, exploring the combination of these two disorders.
Datasets. Given the sensitivity of this topic, large datasets for training deep learning models are not readily available, posing a significant limitation in the field. Existing datasets vary in format and encompass diverse types of information. For instance, some include multiple cognitive tasks, such as image description or fluency tasks. While the majority of datasets focus on Alzheimer’s disease, there are also datasets targeting other conditions, such as MCI.
Most of these datasets are based on audio and text modalities, largely due to privacy concerns associated with video data, particularly when faces are recorded. However, the datasets are relatively small for deep learning purposes, typically containing between 250 and 500 samples, with the largest containing only around 600 samples (excluding DementiaBank, where an exact sample count is difficult to ascertain due to the various corpora it comprises). In terms of language, English is the most commonly represented, present in all but two of the datasets, and in many cases, it is the only language available.
In order to address the issue of data scarcity, various data augmentation techniques can be applied, depending on the specific type of data. In this study, several works that utilize such techniques have been highlighted, showing significant performance improvements compared to not using them. These approaches hold promise for future research directions. An alternative approach to mitigating this issue involves the use of large pre-trained models, particularly in few-shot and zero-shot learning scenarios. These architectures reduce the reliance on large, densely labeled datasets for training downstream tasks. A notable example is the use of Vision Language Models (VLMs) for video understanding [295, 296, 297].
Unimodal Approaches. There are several methods for estimating cognitive states using various information sources. In this study, the focus has been on non-intrusive approaches, including audio, text, images, and video. The most common and relevant modalities are audio and text. Audio data can be analyzed through various deep learning techniques, providing valuable insights into speech articulation and pronunciation. However, one limitation of this modality is the extensive preprocessing required to handle the data effectively. Among these methodologies, Transformer models stand out, with the Wav2Vec2 model, published in 2020, showing the most notable results in this field. While Wav2Vec2 has proven highly effective, more recent Transformer architectures may offer even better performance for audio-based tasks. Investigating how these newer models can be applied to audio data presents an exciting direction for future research.
For textual data, Transformer architectures are the most widely used approach. When combined with linguistic features, they achieve remarkable results. Many studies demonstrate that this architecture outperforms audio-based approaches across different datasets, with few exceptions. This suggests that transcribed text often provides more relevant information, offering deeper insights into how individuals structure their thoughts in sentences, as opposed to how they vocalize them. Conversely, modalities such as image or video are not present in the selected studies due to the scarcity of available datasets. Nonetheless, these modalities can be highly beneficial for estimating cognitive decline, as evidenced by studies analyzing facial expressions during speech or gait patterns in diseases such as Parkinson’s.
Multimodal Approaches. The integration of different modalities has demonstrated significant potential and frequently results in enhanced performance. By integrating various data sources, a more comprehensive understanding of the problem can be achieved. The advantages of a multimodal approach are evident: it facilitates improved insights and fosters the development of more robust solutions. This robustness arises from the model’s capacity to utilize diverse sources of information, reducing its reliance on a single data type for making predictions. Such an approach is particularly beneficial in scenarios where certain modalities may be missing or irrelevant.
However, challenges persist. Processing different types of information necessitates a more complex architecture for effectively combining the modalities. Moreover, the training and inference procedures become more time-consuming due to the requirement for integrating diverse data sources. The application of multimodal strategies, which leverage all available information for a given task, has been shown to enhance the outcomes of unimodal models in 75% of the approaches studied. Conversely, the combination of modalities resulted in inferior performance in 12.5% of the cases, while another 12.5% of studies did not report the performance of unimodal models, making it impossible to conclude that the combination consistently yields superior performance.
This integration can be achieved through the use of pre-existing multimodal models or by fusing modalities. The best results are typically obtained through joint fusion strategies, which generally outperform other fusion techniques in comparative studies. However, the use of pre-existing multimodal models is limited, with only one instance identified in the literature. This indicates a promising avenue for research in this area, particularly concerning multimodal models that incorporate extensive pretraining and rich knowledge bases, which may exhibit substantial performance in this task.
Future directions. One crucial aspect that must be considered, though often overlooked in related works, is the explainability of the models. When dealing with sensitive medical data, it is essential to understand why and how predictions are made in order to extract meaningful insights. Additionally, this information can bring useful insights about the processed data. While recent research has increasingly emphasized explainability, further efforts are needed in this area rather than simply deploying black-box models [298, 299, 300].
@@
An additional potential source of valuable information comes from the use of Human Digital Twins (HDTs). These digital representations of individuals can include comprehensive medical records, offering crucial insights into a patient’s health status. Furthermore, HDTs can simulate and analyze daily human activities, generating synthetic data that could be beneficial for training models aimed at cognitive decline estimation [301, 302].
@@ VIII Conclusions
Cognitive abilities are essential in our daily lives, influencing how we interact with our environment and others. A decline in these abilities over time, as we age, is a normal and natural process. However, in some cases, this deterioration is more severe in some individuals. This severity is often due to different disorders. Early detection of significant cognitive decline can be highly beneficial, allowing for timely professional intervention and improving quality of life.
In this work, we have systematically reviewed the most relevant methodologies for analyzing and predicting this decline, focusing on non-intrusive methods. These approaches do not disrupt daily activities, preserving quality of life without relying on invasive techniques like medical imaging.
To the best of our knowledge, this study represents the first survey of estimation of general cognitive decline based on non-invasive data. The methodologies covered include audio, textual, and visual analyses, leveraging deep learning techniques. We also assess the key features and benefits of each approach, their integration into multimodal frameworks, and their performance on standard datasets. Finally, we present the most significant datasets relevant to this task, providing detailed information about their content, including the activities performed by subjects, the disorders studied, and the modalities provided.
We conclude that there is a significant lack of sufficient datasets in this area, with limited options that are not optimal in terms of sample size. Existing datasets typically focus on audio and text modalities, primarily due to the sensitivity of the data involved. For example, video data can raise additional concerns regarding privacy and ethical considerations, which further limits its availability and use. Consequently, most proposed studies rely on audio and text modalities, while other non-intrusive modalities, including video, have very few representations and often use their own recorded datasets.
@@ Acknowledgments
This work has been funded by the Valencian regional government CIAICO/2022/132 Consolidated group project AI4Health, by Valencian government and International Center for Aging Research ICAR funded project “IASISTEM”. This work has also been supported by a Spanish national and a regional grant for PhD studies, FPU21/00414, and CIACIF/2022/175.

### v2

@@ VII Discussion
VII-A Cognitive disorders
In addition to the natural decline associated with aging, several neurological and neurodegenerative conditions can significantly impair cognitive abilities. The severity and nature of these impairments vary depending on the disorder, and in many cases, they can lead to progressive functional decline and increased dependency. Early detection is crucial for timely intervention that can significantly improve a patient’s quality of life.
Cognitive disorders include a range of conditions such as dementia, MCI, aphasia, Parkinson’s disease, and apathy. As mentioned earlier, the global population is aging, which has led to a growing prevalence of age-related conditions, most notably dementia. As a result, dementia remains the primary focus of research in this area, in part due to its public health impact and the relatively higher availability of relevant datasets. However, there is a growing body of work exploring other disorders, such as MCI, and, in some cases, examining their combined or differential diagnosis.
Notably, the selection of appropriate data modalities for cognitive assessment tasks is inherently task and disorder-specific. This selection should be determined by the nature of the impairments associated with each condition. For instance, Alzheimer’s disease and MCI are primarily characterized by cognitive and linguistic deficits. Patients frequently experience challenges with word retrieval, semantic coherence, and syntactic complexity. These impairments are often detectable in the structure and content of their spoken or written language, making text-based analysis a particularly effective modality for early detection [54, 74].
In contrast, conditions such as Parkinson’s disease and aphasia more directly affect motor control and speech production. These disorders often lead to disruptions in articulation, prosody, phonation, and overall speech fluency. Consequently, audio-based approaches, which can capture features such as jitter, shimmer, intonation, and speech rate, are more suitable for identifying relevant patterns in these populations [299, 300].
This alignment between disorder characteristics and modality relevance is reflected in the literature. Alzheimer’s and MCI are frequently studied using language-based models, whereas Parkinson’s disease and aphasia are more commonly addressed using speech-based models that emphasize vocal features. Understanding these distinctions is essential not only for designing effective predictive models but also for ensuring that research is clinically grounded and tailored to the specific diagnostic needs of each condition.
VII-B Datasets
Due to the sensitivity of this domain, large-scale datasets for training deep learning models remain scarce, posing a significant barrier to advancement. Deep learning techniques typically require extensive and diverse datasets to achieve robust and generalizable performance. This scarcity not only affects model accuracy on benchmark test sets but, more critically, limits the ability of models to generalize beyond their training environments.
Among the limited datasets that are available, there is considerable variation in both structure and content. Some assess performance across multiple cognitive tasks, such as image description or verbal fluency exercises, providing a multifaceted perspective on cognitive function. While most datasets focus on Alzheimer’s disease, some also include data related to conditions such as MCI, thereby expanding their clinical relevance. In terms of data modalities, the majority of resources rely on audio and text. This preference is primarily driven by privacy concerns surrounding video data, mainly when it involves identifiable facial recordings.
From a demographic standpoint, these datasets primarily feature older adults, reflecting the age-related nature of dementia. While this focus aligns with the target population, it reduces the generalizability of models to younger individuals or atypical cases. Gender distribution tends to be relatively balanced. Many existing approaches have been trained on relatively small and homogeneous datasets collected under tightly controlled conditions. While such standardization can enhance internal validity, it often compromises external validity. As a result, models trained under narrow parameters may underperform in real-world environments that differ from the training setup. These models are particularly vulnerable to overfitting and may struggle with variability in speaker demographics, contextual differences, or environmental noise. These imbalances can be addressed through more inclusive data collection practices, as well as through domain adaptation techniques. These methods enable models trained on specific populations to adapt to new target groups, such as individuals from different age brackets, genders, or cultural backgrounds. In addition, the inclusion of relevant metadata, when ethically and legally permissible, allows for more granular performance evaluation and targeted optimization.
Furthermore, the dominance of English in publicly available datasets aggravates this issue. Resources in other languages, such as Spanish or Mandarin, are limited and often lack sufficient diversity and size for reasonable training. This linguistic imbalance can lead to models that underperform on underrepresented populations, reinforcing existing disparities. In clinical domains where fairness, accuracy, and equity are paramount, such biases raise significant ethical concerns. If left unaddressed, these issues could result in misdiagnoses or unequal access to care. This dominance of English in existing datasets can be partially addressed through the use of multilingual and cross-lingual learning strategies. Models pretrained on large-scale multilingual corpora can be fine-tuned using limited data from underrepresented languages, enabling effective knowledge transfer and reducing performance disparities.
Some initiatives have sought to mitigate these biases. For instance, the Taukadial dataset includes both English and Chinese speakers. However, inconsistencies in task design, such as using different pictures for each language group, introduce confounding factors that complicate cross-linguistic comparisons. These protocol variations make it difficult to determine whether differences in model behavior are due to linguistic characteristics or task-related discrepancies.
Fairness-aware machine learning techniques can also be employed during training or post-processing to reduce disparities in model performance across demographic or linguistic subgroups. Approaches such as loss function reweighting, adversarial debiasing, and subgroup-specific threshold calibration can promote more equitable outcomes without significantly compromising overall accuracy. These considerations are particularly critical in healthcare settings, where fairness and transparency are essential to avoid potential harm, misdiagnosis, or unequal access to care.
Despite meaningful technical progress, the clinical adoption of these models remains limited. Key obstacles include the need for large-scale validation across diverse populations, integration with existing clinical workflows, and compliance with regulatory and ethical standards. Moreover, computational limitations are common in clinical settings, where available infrastructure is often optimized for administrative tasks rather than resource-intensive machine learning operations. Effective deployment will require not only more powerful computing systems but also standardized data collection protocols and recording environments aligned with those used during model development.
These limitations underscore the importance of model interpretability. Explainable models can help identify when decisions are due to speech patterns or biased linguistic features. Transparent and accountable decision-making is crucial to ensure fairness, trust, and efficacy, particularly when working with underrepresented or demographically uneven populations.
VII-C Explainability
An essential yet often underemphasized aspect in the development of machine learning models for cognitive assessment is explainability. In the context of sensitive medical data, it is essential to understand why and how predictions are obtained. Explainability not only enhances trust and transparency but also facilitates the extraction of clinically meaningful insights from model behavior. Additionally, this information can bring valuable insights about the processed data. Although recent studies have increasingly acknowledged the importance of explainability, this area remains underdeveloped, with many models still operating as opaque “black boxes” [301, 302, 303].
Among the studies included in this review, explainability methods were notably limited. Nevertheless, some efforts have attempted to address this gap. For instance, authors in [90] examined attention mechanisms within their models to assess the contribution of individual words to final classification outcomes. Their analysis revealed that certain discourse markers, such as “Okay” and “Well” were consistently emphasized by the model while predicting dementia. These markers are normally employed by individuals experiencing word-finding difficulties or requiring additional time to organize their thoughts, providing both interpretability and clinical relevance.
Similarly, Ilias et al. [189] analyzed linguistic patterns using the LIME framework [304] to identify influential tokens, including interjections, personal pronouns, and past-tense verbs. Their findings indicate that individuals with dementia often rely on vague or generalized language, consistent with cognitive impairments. Frequent use of fillers such as “yeah” or “oh” emerged as strong indicators of dementia, reflecting hesitation or lexical retrieval difficulties. These insights not only support the model’s predictions but also align with established clinical observations [305].
Beyond performance metrics, these examples underscore the growing recognition of explainability as a foundational component in the development of trustworthy AI systems for cognitive assessment. While tools such as attention analysis and LIME offer promising avenues for interpretability, their adoption remains limited across the field. Future research should prioritize transparency as a core design principle, embedding explainability throughout the model development, evaluation, and validation pipeline. Establishing standardized protocols to assess and benchmark model explanations will be critical to ensuring clinical reliability, ethical alignment, and equitable deployment across diverse populations.
VII-D Unimodal Approaches
As previously discussed, the clinical profile of each cognitive disorder plays a key role in determining the most informative data modality for assessment. This section focuses on how individual modalities, specifically audio, text, and, to a lesser extent, visual data, have been employed in unimodal machine learning models to estimate cognitive states.
Audio-based approaches rely on speech recordings, which can reveal subtle markers of cognitive decline through articulation, prosody, fluency, and hesitation patterns. Deep learning techniques, particularly Transformer-based architectures, have enabled automatic feature extraction and improved predictive performance in this domain. One of the most widely used models is Wav2Vec2, which learns meaningful acoustic representations from raw audio. However, working with audio data often requires extensive preprocessing, including segmentation, normalization, and noise reduction, all of which add complexity to the modeling pipeline. Moreover, variability in recording conditions and speaker characteristics can impact model generalizability. While Wav2Vec2 remains a strong baseline, emerging models such as HuBERT and Whisper offer promising alternatives that warrant further exploration in cognitive assessment contexts.
Text-based approaches typically analyze transcriptions derived from spoken language. These offer access to syntactic structure, semantic content, and discourse coherence, all of which are often disrupted in individuals with cognitive impairments. Transformer models such as BERT and its derivatives (e.g., DistilBERT, RoBERTa) are widely used in this domain. When combined with handcrafted linguistic features, these models have often outperformed audio-based approaches, highlighting the strength of language structure as a signal of cognitive decline. Notably, most current work employs relatively compact pretrained models (e.g., BERT-base) rather than larger foundation models like LLaMA or Mistral. This trend likely reflects the small size of available datasets and concerns about overfitting with high-capacity models. In some cases, large language models are used not as primary encoders but as feature extractors to enrich downstream representations. Exploring scalable and efficient integration of these larger models remains an open and promising research direction.
Visual modalities, such as image and video data, are underrepresented in the reviewed studies, primarily due to the limited availability of suitable datasets. Nonetheless, these modalities hold substantial potential for cognitive assessment, particularly through the analysis of facial expressions, gaze behavior, and motor features during speech. Prior work in related areas, such as emotion recognition or gait analysis for Parkinson’s disease, demonstrates that visual cues can offer valuable insight into neurological function. However, the limited availability of visual datasets is largely due to the privacy risks associated with video recordings, which frequently contain identifiable personal information. These privacy concerns introduce ethical and legal barriers to data collection and sharing. Potential solutions include data anonymization techniques, such as facial blurring or skeletal pose extraction, which can preserve behavioral features while protecting individual identities. As privacy-preserving technologies continue to evolve, the integration of visual data into cognitive assessment research may become more viable and impactful.
While unimodal approaches have demonstrated promising results, especially when aligned with disorder-specific impairments, they also face limitations in terms of robustness and generalizability. Future work may benefit from combining complementary modalities to capture a richer and more holistic picture of cognitive function.
VII-E Multimodal Approaches
The integration of multiple modalities in cognitive assessment tasks has shown considerable promise, often leading to improved performance and enhanced model robustness. By combining complementary sources of information, such as speech, language, and visual cues, multimodal approaches enable a more comprehensive representation of an individual’s cognitive state. This strategy is particularly beneficial in real-world scenarios where specific modalities may be missing, noisy, or less informative, allowing the model to rely on alternative sources.
Despite these advantages, several challenges persist. Multimodal processing requires more complex model architectures capable of aligning and integrating heterogeneous data types. The training and inference stages are often more computationally intensive, and ensuring meaningful interactions between modalities requires careful design choices. Empirical findings from the reviewed literature reflect both the potential and the limitations of this approach. Specifically, multimodal integration resulted in improved outcomes in 75% of the analyzed studies. However, in 12.5% of cases, performance decreased when combining modalities. While the other 12.5% of studies did not report unimodal baselines, making it difficult to draw definitive conclusions about the universal benefit of multimodality.
Multimodal integration is typically achieved through two main strategies: the use of pretrained multimodal models or the application of fusion techniques. While fusion-based approaches dominate current research, the use of pretrained multimodal models remains limited. Only one notable study was identified in this survey. This gap highlights a promising direction for future work, particularly in exploring models trained on large-scale datasets that encode structured knowledge across modalities. These models have the potential to enhance cognitive assessment by capturing subtle cross-modal interactions and context-dependent behavioral patterns.
In parallel, LLMs, pretrained on vast textual corpora, have demonstrated exceptional performance across a variety of tasks. Many state-of-the-art text-based approaches in this field rely on features derived from these models, underscoring the effectiveness of transfer learning. Extending this paradigm to multimodal settings, pretrained models capable of jointly encoding text, audio, and video inputs represent a compelling opportunity. These architectures can learn integrated multimodal representations that capture complex relationships between language, speech, and behavior, potentially improving accuracy in detecting cognitive decline and related conditions.
Among the fusion strategies examined, joint fusion consistently emerged as the most effective. Early fusion, which combines modalities at the input level, often fails to preserve the distinct characteristics of each data type. Late fusion, which merges predictions at the decision level, may overlook meaningful intermodal relationships. In contrast, joint fusion integrates intermediate feature representations, enabling both modality-specific learning and intermodal interaction modeling. This balanced integration typically leads to superior performance.
Attention-based mechanisms, particularly those implemented within Transformer architectures, have also been explored for cross-modal integration. However, they have not consistently outperformed joint fusion strategies. One possible explanation lies in the preprocessing differences between modalities. Audio data is often segmented into temporal windows, while textual data is tokenized into subword units. These divergent representations may hinder effective alignment in attention layers, limiting the model’s ability to capture fine-grained cross-modal dependencies [306].
Emerging techniques, such as dynamic fusion and hierarchical fusion, offer promising alternatives. These methods enable models to adaptively weigh the contribution of each modality based on task-specific or context-dependent relevance. Such flexibility may further enhance performance and applicability, particularly in clinical settings where input quality and availability can vary significantly.
VII-F Current Limitations
Despite the promising advancements outlined in previous sections, particularly in the use of multimodal architectures and pretrained models, several critical limitations continue to constrain progress in real-world applications of the cognitive state assessment.
One of the most persistent challenges is the scarcity of labeled data, which limits the model’s robustness and generalizability. Multiple reviewed studies have conclusively demonstrated the effectiveness of data augmentation techniques tailored to specific modalities, reporting improved performance compared to models trained without them. These strategies offer a promising avenue for mitigating dataset limitations. Alternatively, the use of large pretrained models, particularly in few-shot and zero-shot learning contexts, can reduce the dependency on densely annotated data. Notably, Vision-Language Models (VLMs) have shown potential for video understanding tasks [307, 308, 281].
Another key limitation lies in the gap between academic benchmarks and real-world clinical implementation. Many studies focus primarily on maximizing performance on curated datasets while overlooking generalizability, usability, and integration into healthcare workflows. Moreover, the computational demands of training and deploying state-of-the-art models pose practical constraints. Video and speech modalities, in particular, require substantial preprocessing and processing power. This is especially problematic in clinical settings where hardware is often optimized for administrative use rather than high-throughput machine learning.
Finally, several overarching issues continue to limit the adoption and trustworthiness of these models. These include overfitting due to small sample sizes, poor generalization across diverse populations and languages, and the lack of interpretability in many high-performing models. Additionally, ethical and regulatory concerns, including transparency, data privacy, and algorithmic fairness, present non-trivial obstacles to clinical deployment. Collectively, these limitations underscore the need for more robust, interpretable, and resource-efficient models, as well as rigorous validation in real-world, heterogeneous environments.
VII-G Future directions
One of the most important directions for future research is the development of large-scale, diverse, and multimodal datasets for training deep learning models in cognitive assessment. These datasets should incorporate detailed demographic information and richer data sources, such as longitudinal medical records and symptom progression data, to enable more personalized and temporally-aware modeling. Including nonverbal cues, such as facial expressions and gestures, may further enhance multimodal approaches. To improve generalizability and fairness, future datasets should also prioritize multilingual and multicultural representation through standardized data collection protocols.
@@ VII-G Future directions
An additional promising source of data comes from the use of Human Digital Twins (HDTs). These digital representations of individuals can include comprehensive medical records, offering crucial insights into a patient’s health status. Furthermore, HDTs can simulate and analyze daily human activities, generating synthetic data that could be beneficial for training models aimed at cognitive decline estimation [309, 310].
@@ VIII Conclusions
This survey reviewed key non-intrusive methodologies for analyzing and detecting this decline, focusing on non-intrusive methods. These approaches do not disrupt daily activities, preserving quality of life without relying on invasive techniques like medical imaging. By bridging technical methods with real-world medical needs, we aim to foster interdisciplinary collaboration on the early detection of cognitive decline.
To the best of our knowledge, this study represents the first survey of estimation of general cognitive decline based on non-invasive data. The methodologies covered include audio, textual, and visual analyses, leveraging deep learning techniques. We also assess the key features and benefits of each approach, their integration into multimodal frameworks, and their performance on standard datasets. We present the most significant datasets relevant to this task, providing detailed information about their content, including the activities performed by subjects, the disorders studied, and the modalities provided.
We conclude that there is a significant lack of datasets in this area, with limited options that are not optimal in terms of sample size. Existing datasets typically focus on audio and text modalities, primarily due to the sensitivity of the data involved. For example, video data can raise additional concerns regarding privacy and ethical considerations, which further limits its availability and use. Consequently, most proposed studies rely on audio and text modalities, while other non-intrusive modalities, including video, have very few representations and often use their own recorded datasets.
@@ Acknowledgments
This work was supported by the Spanish State Research Agency (AEI) under grant: GEMELIA PID2024-161711OB-I00; and ERDF/EU. Additionally, it received support from the CIAICO/2022/132 Consolidated group project “AI4Health”,funded by the Valencian government and International Center for Aging Research ICAR funded project “IASISTEM”. This work is also a part of the ENIA Chair of Artificial Intelligence from the University of Alicante (TSI-100927-2023-6) funded by the Recovery, Transformation and Resilience Plan from the European Union Next Generation through the Ministry for Digital Transformation and the Civil Service. This work has also been supported by a Spanish national and a regional grant for PhD studies, FPU21/00414, and CIACIF/2022/175.